OpenStudy (anonymous):

csc(6x)+cot(6x) / csc(6x)-cot(6x) = _______________. A. cot^2(6x) B. cot(3x) C. cot^2(3x) D. tan(6x) ***my answer: C is that right? @terenzreignz ? :)

OpenStudy (anonymous):

csc(6x)−cot(x)csc(6x)+cot(6x)

terenzreignz (terenzreignz):

methinks you forgot the squares

OpenStudy (anonymous):

tan^2(3x)

OpenStudy (anonymous):

D my friend

OpenStudy (anonymous):

aw darn :( how'd u get that?

OpenStudy (anonymous):

Source(s): MY BRAIN

OpenStudy (anonymous):

lol okay :P thanks!! :D yeah @terenzreignz me thinks so too now :P that stuff confuses me lol

OpenStudy (anonymous):

here ill show my work

OpenStudy (anonymous):

thank you!!! :D

OpenStudy (anonymous):

y = 6x ƶ = y ⁄ 2 = 3x ----------> 2cos²(A ⁄ 2) = 1 + cosA (identity used) ----------> 2sin²(A ⁄ 2) = 1 − cosA (identity used) It is hard to tell what the problem is without brackets [ ] .... [ (1 ⁄ siny) + (cosy ⁄ siny) ] ............ (1 + cosy) = ———————————— ... = ... —————— .... [ (1 ⁄ siny) − (cosy ⁄ siny) ] ............ (1 − cosy) ..... 2cos²(ƶ) = —————— = cot²(ƶ) = cot²(3x)

OpenStudy (anonymous):

wait but didn't you say it was tan(6x) ? :/

OpenStudy (anonymous):

pellet sry stupid typing

OpenStudy (anonymous):

I still haven't mastered my keybord used to windows just switched to apple

OpenStudy (anonymous):

so yes that's what I ment sry

OpenStudy (anonymous):

lol :) ohhh haha cool! are windows and apple that different? pellet? :/

OpenStudy (anonymous):

wait, so is it C or D then? :/

OpenStudy (anonymous):

when you say crap with the s word it switches to pellet

OpenStudy (anonymous):

oh does it really? lol thats funny :P i never knew that :P

OpenStudy (anonymous):

it looks to me like its c

OpenStudy (anonymous):

u agree?

OpenStudy (anonymous):

yeah apple has a completely different keyboard

OpenStudy (anonymous):

yeah, i thought it was C, but you said it was D? ;/ but what you wrote makes sense :) so i guess it's C after all then? :)

OpenStudy (anonymous):

yes ure correct

OpenStudy (anonymous):

haha kk thank you :)

OpenStudy (anonymous):

as I said you have a bright future @iheartfood

OpenStudy (anonymous):

haha thank you @shelbygt520 :)

OpenStudy (anonymous):

Develop the numerator of the expression y: 1/sin 6x+ cot 6x/sin 6x = (1 + cos 6x)/sin 6x Denominator: 1/sin 6x - cos 6x/sin 6x = (1 - cos 6x)/sin 6x. Simplify by sin 6x: y = (1 + cos 6x)/(1 - cos 6x) Use the trig identities: (1+ cos 2a) = 2*cos^2 a, and (1 - sin 2a) = 2sin^2 a Finally: y = 2*cos^2 (3x)/2*sin^2 (3x) = cot^2 (3x).